# Morphology-Specific Peptide Discovery via Masked Conditional Generative Modeling

### v2

@@ Results
where $\odot$ is the element-wise product and $\phi(\cdot)$ is a multilayer perceptron. This summary parameterizes a masked conditional prior $p(z\mid s)$ over the latent variable $z$. Intuitively, users fill only the entries they care about; unspecified fields $i$ are masked ($m_{i}=0$) and are implicitly marginalized by the conditional prior. The decoder then generates the peptide sequence $y$ from $p(y\mid z,s)$. We implement $p$ with an autoregressive Transformer whose cross-attention ”memory” contains a latent token (from $z$) and a condition token (from $s$), allowing the model to honor provided constraints while preserving variability (see Methods). Our approach mirrors the VAEAC’s recipe, conditioning the prior and decoder on the observed context, but uses a learned summary token $s$ instead of injecting raw $c\odot m,m]$ directly.
During training, we apply stochastic masking on top of naturally missing descriptors so the model sees arbitrary subsets of observed fields. For each dataset sample, we randomly form a mask $m$, compute $s$ using Eq. 1, and optimize a CVAE objective that encourages the encoder’s posterior distribution $q(z\mid y)$ to match the masked conditional prior $p(z\mid s)$ while simultaneously maximizing the likelihood of reconstructing the sequence under the decoder $p(y\mid z,s)$. To ensure that $s$ faithfully encodes the provided context (and only the provided context), we add light auxiliary reconstruction terms: a mask-reconstruction head supervises $m$ with a binary cross-entropy (BCE) loss, two heads supervise the binary descriptors with BCE evaluated only where $m_{i}=1$, and a small regressor supervises continuous descriptors with a masked mean-squared error (MSE) loss (again only on unmasked dimensions). These terms regularize $\phi$ to capture the observed specifications without imputing missing entries, and empirically improve constraint satisfaction while mitigating posterior collapse. The full loss formulation is shown in the Methods section.
@@
Novelty, Diversity and Condition Matching
@@
After training, we can discard the encoder component for sampling and leverage the trained mapping of the partial condition sets to the prior distribution and the decoder component as an autoregressive generator. While the generator alone cannot certify assembly or morphology, we can rigorously evaluate (i) novelty, measured relative to the training set, (ii) diversity and the similarity structure of the generated set (to verify guidance without collapse), and (iii) condition matching against the requested descriptors.
@@
To this end, we must generate several peptides with plausible condition sets. Sampling entirely random descriptor combinations is, however, a poor validation strategy, as arbitrary tuples of values (e.g., conditioning on having $\beta$-sheet content with sequence length of four) often lie far off the empirical manifold and are either physically inconsistent or statistically implausible, rendering them effectively impossible targets for PepMorph. Instead, we fit a Gaussian Mixture Model per length ($L\in\{5,\dots,10\}$ for comparison with Njirjak et al. [21]) on the training descriptor space and sample condition vectors from these Gaussian Mixture Models, respecting observed correlations and yielding realistic conditions. Because our model supports partial conditioning, we explicitly mask subsets of descriptors at generation time: for each length $L$, we generate 20 conditions, distributing them evenly across $k\in\{1,\dots,6\}$ used descriptors. We also evaluate 10 rare conditions, randomly split across lengths: five with positive $\beta$-sheet content and five with rare hydrophobic-moment extremes (>0.8>\!0.8), with all other descriptors set near per-length medians. For each of the $130$ conditions, we decode $100$ peptides autoregressively.
@@
We quantify novelty in two ways: as exact sequence matching (fraction of generated sequences present in the training set), and as the nearest-neighbor normalized edit distance (NED) (Levenshtein distance divided by the longer length) to the closest training peptide. The PepMorph model generated highly novel sequences (only $\sim 300$ of the 13k generated sequences are present in the training dataset) that require, on average, roughly a third of their sequence to be edited to match the closest neighbor on the training set ($0.375$ NED, Figure 2c). Additionally, the generated peptides’ amino acid frequency distribution still closely follows the training set’s, and shows no collapse when sampling for rarer conditions (Figure 2b).
@@
As for diversity, we assess it with exact sequence matching within the 100 peptides for each generated condition — showing how well the model explores the space for the same condition set —, but also from the distribution of pairwise NED within each condition (all pairs among the 100 samples) and all conditions (100k random pairs of the 13k generated peptides). The results show that the diversity, like the novelty, is also very high (Figure 2f): generated peptides are considerably different from each other, evidenced by very high pairwise NED but also extremely low exact sequence matching ($0.02\%$), which indicates that the exploration of the search space can be comprehensive.
@@
To characterize similarity among sequences in a way that is sensitive to conditional structure (and not washed out by an extensive training set), we compute the Needleman–Wunsch global alignment with percent identity [21] regarding different sets. We report $\text{Sim}_{\text{train}}$, representing the similarity to all training peptides, $\text{Sim}_{\text{gen}}^{\text{all}}$, representing the similarity to all other generated peptides, and $\text{Sim}_{\text{gen}}^{\text{within}}$, representing the similarity to peptides generated from the same condition only. The results (Figure 2f) reinforce our novelty and diversity claims, as all similarity means are below $10\%$ — comparatively, Njirjak et al.’s genetic algorithm-based model reports $40\%$ similarity across all generated sequences. From these results, two observations follow: first, $\text{Sim}_{\text{train}}$ staying low argues against memorization; second, $\text{Sim}_{\text{gen}}^{\text{within}}$ remains low when compared to $\text{Sim}_{\text{gen}}^{\text{all}}$, implying that partial conditioning preserves substantial sequence variability within each query rather than collapsing to a few templates. The scatter plot in Figure 2e also shows a broad cloud of values: we see no significant difference regarding similarity as the number of conditions increases, but we do see some sets of generated peptides with considerably higher similarity (some reaching close to $20\%$ similarity within conditions), indicating that the type of descriptor that it was conditioned on greatly influences the restriction of the manifold.
@@
Lastly, we quantify condition matching (”effectiveness”, Lim et al. [28]) separately for binary and continuous descriptors. For binary and discrete descriptors, such as SA/no-SA, a sample is a match if the predicted label equals the target. For continuous descriptors, such as AP, a match requires the predicted value to lie within $\pm 10\%$ of the target. We report effectiveness in Figure 2f both per components — (i) length only, (ii) AP only, (iii) SA/no-SA only and (iv) all other descriptors — and as an aggregate score that requires all targeted descriptors to meet their respective criteria simultaneously. This aggregate reflects the model’s ability to honor arbitrary subsets of conditioned descriptors. To be able to evaluate the aggregation description matching, we train a separate AP regressor and self-assembly classifier using the PepMorph aggregation-related subset. As detailed in  Methods, we follow Liu et al. [18]’s approach, but use a single ESM-based transformer backbone [29] with the two task-specific heads: an AP regression head and a self-assembly (SA/no-SA) classification head. This unified model yields state-of-the-art performance on both tasks: the AP head attains an MAE of $0.0393$ (matching Liu et al.’s $0.0391$), and the SA head reaches $96.72\%$ accuracy (vs. their reported $94.49\%$). We predict 3D structures with PEP-FOLD for morphology descriptors and compute the corresponding properties. Overall, condition fidelity is strong ($55\%$ general matching) but descriptor dependent. We see a very high matching rate of length, but considerably lower when measuring against the net charge, $\beta$-sheet content, or hydrophobic moment. One can also see the expected monotonic decline of the all-target rate with increasing $k$ used descriptors in Figure 2d (from $84.58\%$ at $k=1$ down to $24.55\%$ at $k=6$). As more constraints are enforced, any single miss causes failure, and rare settings are intrinsically harder to satisfy. Even so, high length fidelity alongside solid AP/SA matching indicates that the model is not trading constraint satisfaction for trivial length control; rather, PepMorph can juggle multiple design descriptors without collapsing to a single one.
@@
To evaluate our framework against the central hypothesis, we deliberately restrict the conditioning descriptors to the minimum needed to target two morphologies — spherical and fibrillar. As such, we perform an end-to-end quantitative evaluation of morphology control by fixing the condition vector to values intended to generate one of the two target morphologies. These values are duly explained in Table 2. Because the generator only accepts exact descriptor vectors $c$ but peptide space screening requires ranges, we sweep each descriptor over its specified interval with the given step size and evaluate all combinations, collecting the corresponding generated peptides. The only aggregation-related input during generation was the binary aggregation label, which we fixed to $1$ for all conditions.
Table 2: Peptide generation target properties used to direct spherical vs. fibrillar aggregate morphologies. Hydrophobic moment and net charge are shown in their min-max scaled units; targets correspond to high hydrophobic moment and near-neutral net charge. Combinatorial condition grids were sampled with increments of 11 (length), $0.1$ (hydrophobic moment), and $0.05$ (net charge).
@@
Spherical Aggregates Peptide Length 474-7aa Short peptides hardly adopt extended conformations that align into $\beta$-sheets [30].
Hydrophobic Moment 0.61.00.6-1.0 Moderate amphipathicity favors the burial of hydrophobic faces leading to amphiphiles like vesicles or micelles [31].
Net Charge 0.40.60.4-0.6 Near-neutral charge reduces repulsion, enabling self-assembly [26].
Fibrillar Aggregates Peptide Length 7107-10aa Longer peptides may adopt extended conformations that can align into $\beta$-sheets [30].
Has $\beta$-sheet content Yes High strand content drives backbone H-bond stacking into longer fibrils [26].
Net Charge 0.40.60.4-0.6 Near-neutral charge reduces repulsion, enabling self-assembly [26].
@@
First, sequences are screened with the aforementioned AP regressor and SA/no-SA classifier to remove sequences unlikely to aggregate. Next, each candidate peptide is fed into PEP-FOLD to obtain its 3D conformation; from it, together with the sequence, we can recompute all peptide-level descriptors. Only those whose predicted metrics fall within the predefined tolerance ($\pm 10\%$) of the conditioned values are retained. This dual filtering step transforms our pipeline into a reliable end-to-end framework that yields peptides meeting both aggregation and conformation specifications.
Given the defined conditions, we generated $60$ distinct peptides per condition for spherical targets and $300$ for fibrillar targets, yielding $4,800$ peptides for each morphology. After deduplication, the spherical set decreased to $4,697$ sequences, whereas the fibrillar set remained at $4,800$ (no duplicates). We then applied our post-processing filter to prioritize highly aggregating candidates: we retained only peptides with predicted aggregation probability $\geq\!75\%$ from the SA/no-SA classifier, and with predicted aggregation propensity (AP) $\geq\!1.8$ from the regressor. Following this initial screening, $3,986$ spherical and $3,794$ fibrillar sequences remained. After the aggregation-based screening, we assess peptide-level descriptor compliance and predict structures for all remaining sequences with PEP-FOLD, discarding any peptide that fails to satisfy all targeted/conditioned descriptors within the $10\%$ threshold. This yields $64$ spherical candidates and $29$ fibrillar candidates. From the combined candidate set (both morphologies), we select the top $15$ sequences for each morphology by predicted aggregation propensity (AP) from our trained AP regressor, and proceed with validation via CG-MD simulations.
For this particular set of peptides, we can quantitatively distinguish both morphologies from the resulting MD trajectory via the ratio of principal moments of inertia (RMOI) introduced by Wang et al. [23]. For the largest aggregate cluster, let $\lambda_{1}\leq\lambda_{2}\leq\lambda_{3}$ denote the eigenvalues of its inertia tensor (see Methods); then

### v3

@@ 2 Results
where $\odot$ is the element-wise product and $\phi(\cdot)$ is a multilayer perceptron. This summary parameterizes a masked conditional prior $p(z\mid s)$ over the latent variable $z$. Intuitively, users fill only the entries they care about; unspecified fields $i$ are masked ($m_{i}=0$) and are implicitly marginalized by the conditional prior. The decoder then generates the peptide sequence $y$ from $p(y\mid z,s)$. We implement $p$ with an autoregressive Transformer whose cross-attention "memory" contains a latent token (from $z$) and a condition token (from $s$), allowing the model to honor provided constraints while preserving variability (see Methods). Our approach mirrors the VAEAC’s recipe, conditioning the prior and decoder on the observed context, but uses a learned summary token $s$ instead of injecting raw $c\odot m,m]$ directly.
During training, we apply stochastic masking on top of naturally missing descriptors so the model sees arbitrary subsets of observed fields. For each dataset sample, we randomly form a mask $m$, compute $s$ using Eq. 1, and optimize a CVAE objective that encourages the encoder’s posterior distribution $q(z\mid y)$ to match the masked conditional prior $p(z\mid s)$ while simultaneously maximizing the likelihood of reconstructing the sequence under the decoder $p(y\mid z,s)$. To ensure that $s$ faithfully encodes the provided context (and only the provided context), we add light auxiliary reconstruction terms: a mask-reconstruction head supervises $m$ with a binary cross-entropy (BCE) loss, two heads supervise the binary descriptors with BCE evaluated only where $m_{i}=1$, and a small regressor supervises continuous descriptors with a masked mean-squared error (MSE) loss (again only on unmasked dimensions). These terms regularize $\phi$ to capture the observed specifications without imputing missing entries, and empirically improve constraint satisfaction while mitigating posterior collapse. As such, this design accommodates descriptors that are undefined for certain sequences (e.g., PEP-FOLD-derived 3D metrics for very short peptides) and treats them as missing and marginalized rather than imputed. The full loss formulation is shown in the Methods section.
@@
Conditional Generalization and Sample-based Evaluation
@@
After training, we can discard the encoder component for sampling and leverage the trained mapping of the partial condition sets to the prior distribution and the decoder component as an autoregressive generator. While the generator alone cannot certify assembly or morphology, we can rigorously evaluate (i) conditional generalization on held-out test data, (ii) novelty, measured relative to the training set, (iii) diversity and the similarity structure of the generated set (to verify guidance without collapse), and (iv) condition matching against the requested descriptors. Importantly, these metrics are reported as a descriptive properties of the generated ensembles (e.g., to check against memorization or collapse) and are not inherently used as a performance or quality criterion; functional relevance is defined later.
We firstly report a teacher-forced, token-level negative log-likelihood on a held-out test set as a measure for conditional generalization. Concretely, for each test peptide we condition the decoder on the same subset of observed descriptors available in the dataset, as specified by the mask $m$. As seen in Table 2, under this observed-mask conditioning, we obtain a mean per-token cross-entropy of $\approx 0.51$ nats (perplexity $\approx 1.67$). Performance is similar when conditioning on all descriptors for complete-case samples (perplexity $\approx 1.63$) and when applying additional random masking (50%) on top of the observed fields (perplexity $\approx 1.69$). Overall, these results support that the conditional decoder assigns high likelihood to unseen peptides and that its likelihood remains stable under partial conditioning.
Table 2: PepMorph decoder held-out conditional likelihood (teacher-forced). We report the conditional likelihood on the test set under three masking regimes. In the observed-mask case, we evaluate all test samples using all descriptors available for each sample. In the full-mask case, we evaluate the subset of the test set for which all descriptors (aggregation and monomer-level) are available. In the random-on-top case, we use the same subset as the full-mask case but apply additional random masking per sample with probability $p=0.5$ independently for each descriptor. Results are shown as per-token cross-entropy (CE) in nats, with perplexity defined as $\exp(\mathrm{CE})$.
Conditioning mask Per-token CE (nats) Perplexity
Observed mask (all test samples) $\approx 0.51$ $\approx 1.67$
Full mask (complete-case subset) $\approx 0.49$ $\approx 1.63$
Random-on-top (complete-case subset) $\approx 0.52$ $\approx 1.69$
@@
Next, we evaluate sample-based generation quality under plausible condition sets. Sampling entirely random descriptor combinations is, however, a poor validation strategy, as arbitrary tuples of values (e.g., conditioning on having $\beta$-strand assignment with sequence length of five) often lie far off the empirical manifold and are either physically inconsistent or statistically implausible, rendering them effectively impossible targets for PepMorph. Instead, we fit a Gaussian Mixture Model per length ($L\in\{5,\dots,10\}$ for comparison with Njirjak et al. njirjakReshapingDiscoverySelfassembling2024a) on the training descriptor space and sample condition vectors from these Gaussian Mixture Models, respecting observed correlations and yielding realistic conditions. Because our model supports partial conditioning, we explicitly mask subsets of descriptors at generation time: for each length $L$, we generate 20 conditions, distributing them evenly across $k\in\{1,\dots,6\}$ used descriptors. We also evaluate 10 rare conditions, randomly split across lengths: five with positive $\beta$-strand assignment and five with rare hydrophobic-moment extremes (>0.8>\!0.8), with all other descriptors set near per-length medians. For each of the $130$ conditions, we decode $100$ peptides autoregressively.
Table 3: Generation metrics summary. We report novelty via exact sequence matches to training sequences and the median normalized edit distance (NED) to the most similar (nearest-neighbor) sequence in the training dataset. Diversity is reported via exact sequence matches among newly generated sequences and the pairwise NED across all generated sequences. Similarity is computed using Needleman–Wunsch global alignment with percent identity within training samples, between generated and training sequences, and within generated sequences. Condition matching is reported as the percentage of descriptor matches for the generated sequences. Overall, we observe high novelty, broad diversity, low global similarity, and strong condition fidelity.
Metric Value
Novelty
Exact sequence match 2.30%
Nearest-neighbor median NED 0.3750
Diversity
Exact sequence match (mean within conditions) 0.02%
Pairwise mean NED (all conditions) 0.9039
Pairwise std. NED (all conditions) 0.0928
Similarity (NW % id.)
$\mathrm{Sim}_{\text{train}}$ — mean 9.60%
$\mathrm{Sim}_{\text{train}}$ — median 9.59%
$\mathrm{Sim}_{\text{gen}}^{\text{all}}$ — mean 9.89%
$\mathrm{Sim}_{\text{gen}}^{\text{all}}$ — median 9.91%
$\mathrm{Sim}_{\text{gen}}^{\text{within}}$ — mean 9.40%
$\mathrm{Sim}_{\text{gen}}^{\text{within}}$ — median 9.16%
Condition matching
Length 99.72%
AP 81.73%
SA/no-SA 84.34%
Remaining descriptors 61.12%
All 55.21%
@@
We quantify novelty in two ways: as exact sequence matching (fraction of generated sequences present in the training set), and as the nearest-neighbor normalized edit distance (NED) (Levenshtein distance divided by the longer length) to the closest training peptide. The PepMorph model generated highly novel sequences (only $\sim 300$ of the 13k generated sequences are present in the training dataset, $97.7\%$ of them are novel) that require, in median, roughly a third of their sequence to be edited to match the closest neighbor on the training set ($0.375$ NED, Figure 2c). These rare exact matches are expected when sampling from a probabilistic conditional model (and would be far more frequent under memorization). Additionally, the generated peptides’ amino acid frequency distribution still closely follows the training set’s, and shows no collapse when sampling for rarer conditions (Figure 2b).
@@
As for diversity, we assess it with exact sequence matching within the 100 peptides for each generated condition — showing how well the model explores the space for the same condition set —, but also from the distribution of pairwise NED within each condition (all pairs among the 100 samples) and all conditions (100k random pairs of the 13k generated peptides). The results show that the diversity, like the novelty, is also very high (Table 3): generated peptides are considerably different from each other, evidenced by very high pairwise NED but also extremely low exact sequence matching ($0.02\%$), which indicates that the exploration of the search space can be comprehensive.
@@
To characterize similarity among sequences in a way that is sensitive to conditional structure (and not washed out by an extensive training set), we compute the Needleman–Wunsch global alignment with percent identity njirjakReshapingDiscoverySelfassembling2024a regarding different sets. We report $\text{Sim}_{\text{train}}$, representing the similarity to all training peptides, $\text{Sim}_{\text{gen}}^{\text{all}}$, representing the similarity to all other generated peptides, and $\text{Sim}_{\text{gen}}^{\text{within}}$, representing the similarity to peptides generated from the same condition only. The results (Table 3) indicate low global percent identity in our sampled ensembles (all similarity means below $10\%$), consistent with broad sampling under partial conditioning. This showcases that a broader conditional distribution was learned, as the method is not bound to a population-based optimization like Njirjak et al.’s, which yield substantially higher similarity within their genetic algorithm outputs. From these results, two observations follow: first, $\text{Sim}_{\text{train}}$ staying low argues against memorization; second, $\text{Sim}_{\text{gen}}^{\text{within}}$ remains low when compared to $\text{Sim}_{\text{gen}}^{\text{all}}$, implying that partial conditioning preserves substantial sequence variability within each query rather than collapsing to a few templates. The scatter plot in Figure 2e also shows a broad cloud of values: we see no significant difference regarding similarity as the number of conditions increases, but we do see some sets of generated peptides with considerably higher similarity (some reaching close to $20\%$ similarity within conditions), indicating that the type of descriptor that it was conditioned on greatly influences the restriction of the manifold.
@@
Lastly, we quantify condition matching ("effectiveness", Lim et al. limMolecularGenerativeModel2018) separately for binary and continuous descriptors. For binary and discrete descriptors, such as SA/no-SA, a sample is a match if the predicted label equals the target. For continuous descriptors, such as AP, a match requires the predicted value to lie within $\pm 10\%$ of the target. We report effectiveness in (Table 3) both per components — (i) length only, (ii) AP only, (iii) SA/no-SA only and (iv) all other descriptors — and as an aggregate score that requires all targeted descriptors to meet their respective criteria simultaneously. This aggregate reflects the model’s ability to honor arbitrary subsets of conditioned descriptors. To be able to evaluate the aggregation description matching, we train a separate AP regressor and self-assembly classifier using the PepMorph aggregation-related subset. As detailed in  Methods, we follow Liu et al. liuEfficientPredictionPeptide2023a’s approach, but use a single ESM-based transformer backbone linLanguageModelsProtein2022a with the two task-specific heads: an AP regression head and a self-assembly (SA/no-SA) classification head. This unified model yields state-of-the-art performance on both tasks: the AP head attains an MAE of $0.0393$ (matching Liu et al.’s $0.0391$), and the SA head reaches $96.72\%$ accuracy (vs. their reported $94.49\%$).
We predict 3D structures with PEP-FOLD for morphology descriptors and compute the corresponding properties. Overall, condition fidelity is strong ($55\%$ general matching) but descriptor dependent. We see a very high matching rate of length, but considerably lower when measuring against the net charge, $\beta$-strand assignment, or hydrophobic moment. One can also see the monotonic decline of the all-target rate with increasing $k$ used descriptors in Figure 2d (from $84.58\%$ at $k=1$ down to $24.55\%$ at $k=6$). Even so, high length fidelity alongside solid AP/SA matching indicates that the model is not trading constraint satisfaction for trivial length control; rather, PepMorph can juggle multiple design descriptors without collapsing to a single one. This decline, however, highlights an expected trade-off: as the number of simultaneously fixed descriptors $k$ increases, the probability that a sampled peptide satisfies all requested constraints (under our $\pm 10\%$ tolerance for continuous targets) decreases, as any single mismatch causes failure; not only that, but also potentially rarer setting combinations become intrinsically harder to satisfy. In practice, this means PepMorph is most effective when users fix only the few descriptors that are essential to the design intent and leave the remainder unrestrained (via masking), using post-generation screening as a safeguard rather than attempting to "hard-satisfy" many constraints at once.
@@
To evaluate our framework against the central hypothesis, we deliberately restrict the conditioning descriptors to a small set with values within specific ranges that could act as morphology proxies for two regimes — spherical and fibrillar, as in Table 4 —, enriching the probability of obtaining one of the two target morphologies (not uniquely determining them). As such, we perform an end-to-end quantitative evaluation of morphology steering by fixing the condition vector to values in those ranges, but also positively fixing the self assembly descriptor is_assembled to enforce that aggregation is a prerequisite for morphology formation — although not a morphology descriptor. The ranges of the chosen descriptors for each target morphology are duly explained in Table 4. Because the generator only accepts exact descriptor vectors $c$ but peptide space screening requires ranges, we grid-sweep each descriptor over its specified range with a fixed step size and evaluate all combinations, collecting the corresponding generated peptides. The only aggregation-related input during generation was the binary aggregation label, which we fixed to $1$ for all conditions.
Table 4: Target property grids for morphology-directed peptide generation. Hydrophobic moment and net charge are shown in min–max scaled units; targets correspond to high hydrophobic moment and near-neutral net charge. Combinatorial condition grids were sampled with increments of 11 (length), $0.1$ (hydrophobic moment), and $0.05$ (net charge). We treat these ranges as heuristic steering windows rather than deterministic rules, and validate morphology only via downstream CG-MD later on.
@@
Spherical Aggregates Peptide Length 575-7aa Short peptides are less likely to adopt extended conformations that are characteristic of elongated aggregates stangerLengthdependentStabilityStrand2001.
Hydrophobic Moment 0.61.00.6-1.0 Moderate amphipathicity favors the burial of hydrophobic faces leading to amphiphiles like vesicles or micelles familiaPredictionPeptideProtein2015.
Net Charge 0.40.60.4-0.6 Near-neutral charge reduces repulsion, enabling self-assembly dehsorkhiSelfassemblingAmphiphilicPeptides2014.
Fibrillar Aggregates Peptide Length 7107-10aa Longer peptides may adopt extended conformations that can align into $\beta$-sheets stangerLengthdependentStabilityStrand2001.
Has $\beta$-strand Yes High monomeric $\beta$-strand assignment drives backbone H-bond stacking into longer fibrils dehsorkhiSelfassemblingAmphiphilicPeptides2014.
Net Charge 0.40.60.4-0.6 Near-neutral charge reduces repulsion, enabling self-assembly dehsorkhiSelfassemblingAmphiphilicPeptides2014.
@@
First, sequences are screened with the aforementioned AP regressor and SA/no-SA classifier to remove sequences unlikely to aggregate. Next, each candidate peptide is fed into PEP-FOLD to obtain its 3D conformation; from it, together with the sequence, we can recompute all peptide-level descriptors. Only those whose predicted metrics fall within the predefined tolerance ($\pm 10\%$) of the conditioned values are retained. This dual filtering step transforms our pipeline into a reliable end-to-end framework that yields peptides meeting both aggregation and proxy specifications.
To keep a matched sampling budget across the two target regimes, we generated a fixed total of $4,800$ sequences per target. Given that the condition grids have a different number of possible combinations, this corresponded to $80$ distinct peptides per condition for spherical targets and $300$ for fibrillar targets. After deduplication, the spherical set decreased to $4,778$ sequences, whereas the fibrillar set remained at $4,800$ (no duplicates). We then applied our post-processing filter to prioritize highly aggregating candidates: we retained only peptides with predicted aggregation probability $\geq\!75\%$ from the SA/no-SA classifier, and with predicted aggregation propensity (AP) $\geq\!1.8$ from the regressor. Following this initial screening, $4,173$ spherical and $3,794$ fibrillar sequences remained. After the aggregation-based screening, we assess peptide-level descriptor compliance and predict structures for all remaining sequences with PEP-FOLD, discarding any peptide that fails to satisfy all targeted/conditioned descriptors within the $10\%$ threshold. This yields $76$ spherical candidates and $29$ fibrillar candidates. From the combined candidate set (both morphologies), we select the top $15$ sequences for each morphology by predicted aggregation propensity (AP) from our trained AP regressor, and proceed with validation via CG-MD simulations.
@@
To verify whether sequences that match the proxies arise from conditioning during generation rather than post-generation rejection sampling alone, we also evaluated two baselines under the same screening pipeline and the same sampling budget (4,800 sequences per target regime). First, we generated 4,800 uniformly random peptides and applied the full aggregation and descriptor filtering pipeline; this yielded no spherical candidates and only 4 fibrillar candidates. Second, we used PepMorph in a length-only mode by masking all morphology-proxy descriptors at generation time and relying solely on post-generation screening; under the same budget, this yielded only 3 spherical candidates and 7 fibrillar candidates. These results, explicitly shown in Table 5, indicate that sampling without morphology conditioning is markedly sample-inefficient, and that conditioning during generation is essential to enrich the candidate pool before MD validation. Because PepMorph is a sampling-based constrained discovery pipeline rather than an iterative optimization loop, we quantify efficiency by the probability of obtaining a proxy-compliant candidate (or a "proxy-hit") under a finite screening budget, which directly implies the expected number of screening calls per hit. We report corresponding Wilson 95% confidence intervals, calls-per-hit estimates and budget–success curves, as well as the full funnels for both of these baselines, in Supplementary Material.
Table 5: Baselines for morphology-targeted discovery via post-filtering. Under the same sampling budget (N=4,800N=4{,}800 generated sequences per target regime) and the same post-generation screening pipeline, conditioning on morphology-proxy descriptors substantially increases the yield of proxy-compliant candidates. Calls-per-hit are computed as $N/\#\text{hits}$ (infinite if zero hits).
Generator / sampling method Sphere-target Fibril-target
PepMorph model, all proxies 76 / 4,800 ($\approx 63$ calls/hit) 29 / 4,800 ($\approx 166$ calls/hit)
PepMorph model, length-only 3 / 4,800 ($1600$ calls/hit) 7 / 4,800 ($\approx 686$ calls/hit)
Random sequences (uniform residues) 0 / 4,800 ($\infty$ calls/hit) 4 / 4,800 ($1200$ calls/hit)
For the chosen set of peptides for CG-MD validation, we can quantitatively distinguish both morphologies from the resulting MD trajectory via the ratio of principal moments of inertia (RMOI) introduced by Wang et al. wangAggregationRulesShort2024. For the largest aggregate cluster, let $\lambda_{1}\leq\lambda_{2}\leq\lambda_{3}$ denote the eigenvalues of its inertia tensor (see Methods); then